OpenStudy (anonymous):

Simplify (-3+6i)/(-1+i)

OpenStudy (anonymous):

does "simplify" mean "write in standard form as \[a+bi\]? is so, multiply top and bottom by the conjugate of the denominator

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what do u mean by conjugate?

OpenStudy (anonymous):

\[(a+bi)(a-bi)=a^2+b^2\] a real number. so in your case you want \[\frac{-3+6i}{-1+i}\times \frac{-1-i}{-1-i}=\frac{(-3+6i)(-1-i)}{1^2+1^2}\]

OpenStudy (anonymous):

all the work is multiplying in the numerator. the denominator is 2

OpenStudy (anonymous):

do i multiply the top out?

OpenStudy (anonymous):

yes you have to multiply out in the numerator

OpenStudy (anonymous):

when you are all through you should get \[\frac{9-3i}{2}\] which in standard form is \[\frac{9}{2}-\frac{3}{2}i\]